A car starting from rest accelerates for 15 min. until its velocity is 20 m/s. It then moves at a constant velocity from another 20 min. before it slows down and finally stopped in another 10 m. Find the acceleration during the first 15 min.


1
Expert's answer
2021-09-09T10:49:28-0400

Given:

"v_0=0,"

"t_1=15\:\rm min=900\: s"

"v_1=20\:\rm m/s"


The acceleration during the first 15 min

"a=\frac{\Delta v}{\Delta t}=\frac{v_1-v_0}{t_1}"

"a=\frac{20\:\rm m/s-0\: m/s}{900\:\rm s}=0.022\:\rm m/s^2"